What is the hydronium (H3O+) concentration of a solution with a hydroxide (OH-) conccentration of 2.3 × 10-8?What is the hydronium (H30+) concentration of a solution with a hydroxide (OH-) conccentration of 2.3 x 10-8?
Answer #1

Ionic product of water(Kw) at STP is 10^-14

[H3O+][OH-] = Kw

[OH-] = 2.3*10^-8 M

[H3O+]*(2.3*10^-8) = 10^-14

[H+] = (10^-14)/(2.3*10^-8) M

= 4.3*10^-7 M